OpenStudy (anonymous):

The midpoints of the sides of a triangle are A(2,-1), B(1,3), and C(-3,1). Determine the vertices of the triangle.

OpenStudy (anonymous):

Le (a,b), (c,d) and (e,f) be the vertices, then \[ a+e=2\\ b+f=6\\ c+e=4\\ d+f=-2\\ a+c=-6\\ b+d=4 \] Solve them and you find \[\{a=-4,c=-2,b=6,d=-2,e=6,f=0\} \]

OpenStudy (anonymous):

How did you determine the equations? a+e=2 b+f=6 c+e=4 d+f=−2 a+c=−6 b+d=4

OpenStudy (anonymous):

Can you explain it please?

OpenStudy (anonymous):

B=(1,3) is the midpoint of (a,b) and (e,f) then a + e = 2(1) and b+f = 2(3) Same for the others

OpenStudy (anonymous):

oh, I see. I think I understand now. Thank You.

OpenStudy (anonymous):

I got some different numbers for f, b, and d. f=6-b 6-b+d=-2 d= -2-6+b d=-8+b -8+b+b=2 -8+2b=2 2b=10 b=5 _________________________ d+5=2 d=-3 ___________________________- f=6-5 f=1

OpenStudy (anonymous):

I am not sure which parts are incorrect